In a recent survey a researcher claimed that on average filipino teenagers watch k-dramas for 3.5 hours daily is the claim correct if a random sample of 50 filipino teenagers showed a mean of 2.8 hours with a standard deviation of 0.9 hours? use 99% confidence level

1
Expert's answer
2022-06-01T03:30:24-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=3.5"

"H_1:\mu\not=3.5"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=49" and the critical value for a two-tailed test is "t_c =2.679952."

The rejection region for this two-tailed test is "R = \{t:|t|>2.679952\}."


The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{2.8-3.5}{0.9/\sqrt{50}}=-5.4997"


Since it is observed that "|t|=5.4997>2.679952=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=39" degrees of freedom, "t=-3.1623" is "p=0.000001," and since "p= 0.000001<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 3.5, at the "\alpha = 0.01" significance level.